• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name V. Sulkosky

    Log entry time 16:03:58 on April13,2008

    Entry number 231774

    keyword=Day Shift Summary

    April 02, 2008 Swing Shift Summary

    Swing Shift Crew:
    V. Sulkosky (SL), J. Udias (TO), C. Fernandez-Ramirez (3rd)

    Summary
    =======
    08:00 Start shift with production at 2.5 uA on Loop 1 LH2.
    08:58 Noticed that BPMA Y position drifted from 0.7 mm up to 1.6 mm.
    Asked MCC to bring it back to 0.7 mm. This affects runs 2896,
    2897 and 2898.
    09:08 Beam position finally back to nominal positions.
    09:42 Noticed that DT was high 90% and R0C5 was in a bad state during run
    2900. Rebooted ROC5, but ROC3 then had problems. Rebooted all ROC's
    and restarted RunControl.
    10:15 Beam has been tripping due to Hall C at 80 uA for the past 30 mins.
    13:52 Noticed high DT 94% near the end of 2918, ROC5 was rebooted and CODA restarted.
    14:07 BigBite dE, E and WC HV's tripped off. Asked for beam off.
    14:17 BigBite HVs restored and ask for beam back.
    14:22 DAQ failed to start.
    14:25 HV tripped off again.
    14:37 Another HV trip, which are related to the BPMA and BPMB Y positions being
    considerably off after a beam trip.
    14:50 Resumed data taking.
    14:58 Another HV trip.
    15:07 Resumed data taking.

    Left HRS Run Summary
    ====================
    momentum = 1.007 GeV/c, theta = 12.5 degs.



    Start Run Events Comments
    ========================================================================
    08:01 2893 1.0M Production on LH2
    08:13 2894 1.0M Production on LH2
    08:25 2895 1.0M Production on LH2
    08:38 2896 1.0M Production on LH2, BPMA Y shifted
    08:52 2897 1.0M Production on LH2, BPMA Y shifted
    09:04 2898 1.0M Production on LH2, BPMA Y shifted
    09:16 2899 1.0M Production on LH2
    09:29 2900 0.84k Production on LH2, High DT at end of run
    09:54 2902 1.0M Production on LH2
    10:06 2903 1.0M Production on LH2
    10:20 2904 1.0M Production on LH2
    10:33 2905 1.0M Production on LH2
    10:49 2906 1.0M Production on LH2
    11:02 2907 1.0M Production on LH2
    11:16 2908 1.0M Production on LH2
    11:28 2909 1.0M Production on LH2
    11:41 2910 1.0M Production on LH2
    11:52 2911 1.0M Production on LH2
    12:06 2912 1.0M Production on LH2
    12:22 2913 1.0M Production on LH2
    12:36 2914 1.0M Production on LH2
    12:49 2915 1.0M Production on LH2
    13:02 2916 1.0M Production on LH2
    13:21 2917 1.0M Production on LH2
    13:33 2918 0.78k Production on LH2, High DT before the end
    13:58 2919 0.76k Production on LH2, BigBite HVs tripped off.
    14:50 2922 0.57k Production on LH2
    15:07 2923 1.0M Production on LH2
    15:23 2924 1.0M Production on LH2
    15:39 2925 1.0M Production on LH2
    15:57 2926 ???M Production on LH2, continue into swing





    Right HRS Run Summary
    =====================
    momentum = 0.4735 GeV/c, theta = 105.0 deg


    Start Run Events Comments
    ================================================================
    08:01 20665 28.4k corresponds to LHRS runs 2893-2894
    08:26 20666 207k corresponds to LHRS runs 2895-2908
    11:28 20667 204k corresponds to LHRS runs 2909-2919
    14:22 20668 ???k corresponds to LHRS runs 2922-2926, continue into swing